# SGSST: Scaling Gaussian Splatting Style Transfer

### v1

@@ Appendix C ARF and high resolution inputs
Appendix D Details on the perceptual study
As described in the main paper, a comparative perceptual study was conducted using the 40 3D style transfer experiments presented in Section B (Figures 17 to 25). For each experiment, they were asked to pick the image that appeared to be the most faithful to the style image among the three displayed results. Each participant was shown ten random experiments and participation was voluntary. Figure 26 is an example of such an experiment displaying the style image (top left image) and three views of the scenes stylized by SGSST, ARF and StyleGaussian respectively and displayed in random order. To choose between one of the three results, the participant had to press the left arrow key to select the bottom left result, the up arrow key to select the top right result and the down arrow key to select the bottom right result.

### v2

@@ Appendix D ARF and high resolution inputs
Appendix E Details on the perceptual study
As described in the main paper, a comparative perceptual study was conducted using the 40 3D style transfer experiments presented in Section C (Figures 18 to 26). For each experiment, they were asked to pick the image that appeared to be the most faithful to the style image among the three displayed results. Each participant was shown ten random experiments and participation was voluntary. Figure 28 is an example of such an experiment displaying the style image (top left image) and three views of the scenes stylized by SGSST, ARF and StyleGaussian respectively and displayed in random order. To choose between one of the three results, the participant had to press the left arrow key to select the bottom left result, the up arrow key to select the top right result and the down arrow key to select the bottom right result.